OpenStudy (anonymous):

Please help! I will fan+medal The nth term of a geometric sequence is given by an=27(0.1)^n-1 What are the first five terms of the sequence? What is the sum of the first 10 terms of the sequence?

OpenStudy (jdoe0001):

have you covered geometric sequences yet?

OpenStudy (anonymous):

yes I believe the first 5 terms are 27, 2.7, 0.27, 0.027, and 0.0027

OpenStudy (anonymous):

is that correct?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

to get the sum of the first 10 terms we'd use the "sum of a finite geometric sequence" formula thus \(\large \bf a_{\color{brown}{ n}}=27({\color{blue}{ 0.1}})^{{\color{brown}{ n}}-1}\qquad S_{\color{brown}{ n}}=a_1\left(\cfrac{1-{\color{blue}{ r}}^{\color{brown}{ n}}}{1-{\color{blue}{ r}}}\right)\)

OpenStudy (jdoe0001):

so the get the " S "um of the 10 terms just set n = 10

OpenStudy (anonymous):

i got 10.46 as my answer

OpenStudy (anonymous):

is that correct?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\bf \large { a_{\color{brown}{ n}}=27({\color{blue}{ 0.1}})^{{\color{brown}{ n}}-1}\qquad S_{\color{brown}{ n}}=a_1\left(\cfrac{1-{\color{blue}{ r}}^{\color{brown}{ n}}}{1-{\color{blue}{ r}}}\right) \\ \quad \\ S_{\color{brown}{ 10}}=27\left(\cfrac{1-{\color{blue}{ 0.1}}^{\color{brown}{ 10}}}{1-{\color{blue}{ 0.1}}}\right) }\) I seem to get something a bit bigger

OpenStudy (anonymous):

hmm so my answer is 29.9?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

it can easy round to 30.. but yes.. 29.999999997

OpenStudy (anonymous):

can you help me out with a few more questions?

OpenStudy (jdoe0001):

sure... just post anew... if I dunno someone else may know, and we can also revuse each other

OpenStudy (anonymous):

Okay ill close this and post a new one